# Fully ergodic simulations using radial updates

### v1

@@ I Introduction
The Hybrid Monte Carlo (HMC) method is one of the most successful tools in the simulation of lattice field theories. However, despite its many advantages, its application often faces challenges due to manifolds of vanishing fermion determinant which result in infinite potential barriers that lead to diverging force terms in the molecular dynamics evolution. This separates regions in configuration space and results in ergodicity violations of the algorithm. One important example is the freezing of global topological fluctuations in lattice quantum chromodynamics (QCD) discretizations with chiral fermions [1] or near the continuum limit, where the integer-valued topological charge cannot be smoothly changed by molecular dynamics. Another prime example that exhibits configuration space partitioning is the fermionic Hubbard model in the particle/hole basis using the exponential discretization and formulated in terms of a non-compact auxiliary field by means of a continuous Hubbard-Stratonovich transformation. In this scenario, the occurring fermion determinants vanish on manifolds of codimension 1 and therefore trigger an ergodicity problem in HMC simulations [2]. Reliable application of HMC necessitates the development of strategies to circumvent the potential barriers; Ref. [2] explores coarser molecular dynamics integration and discrete jumps while other approaches introduce a complexified auxiliary field over which to integrate [3, 4, 5, 6]. Normalizing flows represent a more radical departure from HMC for sampling these models [7] and do not use molecular dynamics evolution to generate the proposed updates. In this work, we propose another method, which interleaves the HMC simulation with so-called radial updates to overcome the ergodicity problem posed by the infinite potential barriers. Radial updates refer to multiplicative Metropolis-Hastings updates in a radial direction of non-compact fields that enable jumps over potential barriers and thus ensure ergodicity at comparably low computational cost.
@@ II Hybrid Monte Carlo
Critical to the correctness of HMC is that the combination of refreshing the momentum and MD integration must be able to reach every configuration of $\phi$ of nonzero weight. If infinite potential barriers partition the configuration space, the MD trajectory will never reach some valid configurations, leading to an unwanted bias in observable estimates. Notice that when $S$ diverges so too does the force $F[\phi]=-\frac{\partial S}{\partial\phi}$ (6) and the MD trajectory is rightly repelled. But if $S$ diverges everywhere on a codimension-1 manifold, the MD evolution becomes trapped and cannot reach every configuration. This bias is called an ergodicity problem.
In a practical implementation of HMC, the integration of Hamilton’s equation is inevitably done numerically. To this end, a trajectory of length $T$ is discretized into $N_{\mathrm{MD}}$ equal steps of size $\epsilon=T/N_{\mathrm{MD}}$. An initial sample $(\phi,\pi)\equiv(\phi(0),\pi(0))$ is then evolved to a new sample $(\tilde{\phi},\tilde{\pi})\equiv(\phi(T),\pi(T))$ using a numerical integration scheme. While an area-preserving and reversible integrator ensures the phase space is conserved through the integration, discretization errors can introduce violations of the energy conservation. Therefore, to retain detailed balance, the new configuration $\tilde{\phi}$ is proposed to the Markov chain and accepted according to the Metropolis test with probability,
@@ IV.1 Formulation
Here, $h=\Delta_{t}\delta_{\langle z,z^{\prime}\rangle}$ is the hopping matrix and $\mathcal{B}_{t}$ encodes the anti-periodic boundary conditions with $\mathcal{B}_{t}=+1$ for $0<t^{\prime}<N_{t}$ and $\mathcal{B}_{0}=-1$. This exponential discretization of the Hubbard model will serve as the basis for the analysis carried out in the remainder of this work. Other discretizations exist, but the exponential discretization is attractive because it preserves an exact chiral symmetry of the Hubbard Hamiltonian (12). However, that same symmetry implies that the fermion determinant $\det\left(M[\phi|\kappa]M[-\phi|-\kappa]\right)$ has domain walls of zero weight with codimension 1[2], and the action (15) diverges. As discussed in Section II, this makes HMC susceptible to ergodicity problems caused by infinite potential barriers that need to be addressed to ensure correct sampling.
@@ IV.2 Two-site model: restoring ergodicity
which eliminates autocorrelations arising from the harmonic part of the action [21]. The number of MD steps is tuned to obtain a fine integrator with acceptance rate $>99\%$, leading to $N_{\mathrm{MD}}=60$ for $N_{x}=2$, $N_{t}=1$. We begin by generating $10^{4}$ configurations, saving the configuration after each HMC step. In the simulation with radial updates, we additionally employ one radial update per HMC step with standard deviation $\sigma_{R}=1.8$.
@@
Figure 2: HMC configurations (red dots) for the two-site model on a single time slice ($N_{t}=1$) for $U=18$, $\beta=1$ and $\kappa=1$. The configurations are compared to the exact distribution, shown as a contour plot. The HMC simulations were conducted using (17) and $N_{\mathrm{MD}}=60$, achieving an acceptance rate $>99\%$. The left panel shows the simulation without radial updates, while the right panel includes radial updates. In the latter, a single radial update was performed before each HMC step with a proposal standard deviation of $\sigma_{R}=1.8$. Each panel shows $10^{4}$ recorded configurations, with measurements taken after each HMC step. The plots at the margins compare the exact marginal distribution to the histograms obtained from the visualized trajectories.
@@
Next, we increase the dimensionality $d$ of the system by increasing the number of time slices $N_{t}$, while keeping the parameters $U$, $\beta$, and $\kappa$ unchanged. The trajectory length $T$ is adjusted to reduce autocorrelations (17) and $N_{\mathrm{MD}}$ is tuned to maintain a fine integrator with an acceptance rate $>99\%$. The specific parameter choices and further details on the simulations are given in Appendix B.2. In this case, the potential barriers can still be visualized by considering $\Phi_{x}=\sum_{t}\phi_{tx}$, because in the strong-coupling limit $U/\kappa\gg 1$, the probability weights of a configuration $\Phi=(\Phi_{1},\Phi_{2})$ are well-approximated by the exact one-site distribution derived in Ref. [2]. We begin by examining the example case of $N_{t}=8$, visualizing $10^{4}$ recorded configurations, both with and without radial updates. In both simulations, we recorded a configuration only after every 10th step, such that the shown trajectory corresponds to a total of $10^{5}$ iterations of the respective algorithm. In the simulation with radial updates we employ one radial update with standard deviation $\sigma_{R}=0.6$ per HMC step. The results are shown in Fig. 3. Similarly to the $N_{t}=1$ case, we observe that without radial updates, the trajectory remains confined to the middle diagonal band in the two-dimensional $\Phi$-plane. However, with radial updates, the algorithm efficiently explores the entire configuration space, effectively restoring ergodicity.
@@
Figure 3: HMC configurations (red dots) for the two-site model with $N_{t}=8$ for $U=18$, $\beta=1$ and $\kappa=1$. The configurations $\Phi=(\Phi_{1},\Phi_{2})$ with $\Phi_{x}=\sum_{t}\phi_{tx}$ are compared to the exact one-site distribution, recovered in the strong-coupling limit $U/\kappa\gg 1$ and shown as a contour plot. The HMC simulations were conducted using an optimal trajectory length (17) and $N_{\mathrm{MD}}=60$, achieving an acceptance rate $>99\%$. The left panel shows the simulation without radial updates, while the right panel includes radial updates. In the latter, a single radial update was performed before each HMC step with a proposal standard deviation of $\sigma_{R}=0.6$. Each panel shows $10^{4}$ recorded configurations, with measurements taken after every 10th HMC step. The plots on the margins compare the marginal one-site distribution to the histograms obtained from the visualized trajectories.
@@ IV.3 Two-site model: reducing autocorrelation times
Figure 5: Integrated autocorrelation time $\tau_{\mathrm{int}}$ as a function of the proposal standard deviation $\sigma_{R}$ for the (left) observable $\mathcal{O}_{\Phi}$ (23) and the (right) single particle correlator $\mathcal{O}_{C}$ (19). The fits (green line) are obtained using the ansatz (25), with fit results shown in the corresponding legends. These results are used to estimate the optimal proposal standard deviation $\sigma_{R}^{\mathrm{(min)}}$ (grey dashed line), whose value is also provided in the legends. The underlying model is the two-site model with $N_{t}=8$ for $U=18$, $\beta=1$, and $\kappa=1$. The HMC simulations were conducted using an optimal trajectory length (17) and $N_{\mathrm{MD}}=50$, achieving an acceptance rate $>99\%$. Each HMC step was preceded by a single radial update and a total of $N_{\mathrm{conf}}=3\times 10^{5}$ configurations were recorded, with measurements taken after each HMC step.
@@
Figure 6: The left panel shows a double logarithmic plot of the position of the minimum $\sigma_{R}^{(\mathrm{min})}$ as a function of dimensionality $d=N_{x}N_{t}$, along with fits using ansatz (26) to determine the leading-order scaling. The right panel illustrates a double logarithmic plot of the integrated autocorrelation time at the respective minimum, denoted by $\tau_{\mathrm{int}}^{(\mathrm{min})}$, also as a function of the dimensionality $d$. The data in both panels is fitted to the respective ansatz, (26) and (27), to determine the leading-order scaling. The results were obtained using $\mathcal{O}_{\Phi}$ (23) shown as blue circles, and the single particle correlator $\mathcal{O}_{C}$ (19) projected to the negative-parity irrep (20) shown as red squares. Data points shown in low opacity have been excluded from the fitting procedure and fit results are provided in the legends, where $\alpha^{\prime}=\log_{\mathrm{10}}\alpha$. The underlying model is the two-site Hubbard model with varying $N_{t}$ for $U=18$, $\beta=1$, and $\kappa=1$.
@@ IV.4 Perylene
[…]
@@ V Summary and Outlook
Using the Hubbard model as a primary example, we demonstrated that radial updates successfully restore ergodicity by enabling transitions between regions separated by codimension 1 manifolds of vanishing fermion determinant. Furthermore, we examined the scaling behavior of integrated autocorrelation times with the inclusion of radial updates, consistently observing polynomial scaling, which also indicates the absence of ergodicity violations in practice.
@@ B.2 Two-site model
For the two-site model, we performed a series of simulations for an increasing number of time slices ($N_{t}$), utilizing radial updates with varying proposal standard deviations ($\sigma_{R}$). For each choice of $N_{t}$, we employed an optimal HMC trajectory length (17) and tuned the number of MD steps ($N_{\mathrm{MD}}$) to obtain an acceptance rate of $>99\%$. One radial update was applied prior to each HMC step, and measurements were taken after every HMC step. To obtain reliable estimates of the integrated autocorrelation times, the number of recorded configurations ($N_{\mathrm{conf}}$) was increased incrementally for higher dimensionalities. The specific choices of $N_{t}$, the corresponding $N_{\mathrm{MD}}$, and $N_{\mathrm{conf}}$ used in these simulations are summarized in Tab. 2.
@@ Appendix C Four-site model
Figure 12: Integrated autocorrelation time $\tau_{\mathrm{int}}$ as a function of the proposal standard deviation $\sigma_{R}$ for the observables $\mathcal{O}_{\Phi}$ (left) and $\mathcal{O}_{f}$ (right). The observables are defined in (23) and (30), respectively. The fits (green line) are obtained using the ansatz (25), with fit results shown in the corresponding legends. These results are used to estimate the optimal proposal standard deviation $\sigma_{R}^{\mathrm{(min)}}$ (grey dashed line), whose value is also provided in the legends. The underlying model is the four-site Hubbard model on $2\times 2$ square lattice with $N_{t}=8$ for $U=18$, $\beta=1$, and $\kappa=1$. The HMC simulations were conducted using (17) and $N_{\mathrm{MD}}=50$, achieving an acceptance rate $>99\%$. Each HMC step was preceded by a single radial update and a total of $N_{\mathrm{conf}}=3\times 10^{5}$ configurations were recorded, with measurements taken after each HMC step.
Figure 13: The left panel shows a double logarithmic plot of the position of the minimum $\sigma_{R}^{(\mathrm{min})}$ as a function of dimensionality $d=N_{x}N_{t}$, along with fits using ansatz (26) to determine the leading-order scaling. The right panel illustrates a double logarithmic plot of the integrated autocorrelation time at the respective minimum, denoted by $\tau_{\mathrm{int}}^{(\mathrm{min})}$, also as a function of the dimensionality $d$. The data in both panels is fitted to the respective ansatz, (26) and (27), to determine the leading-order scaling. Fit results are provided in the legends, where $\alpha^{\prime}=\log_{\mathrm{10}}\alpha$. The underlying model is the four-site Hubbard model on a $2\times 2$ square lattice with varying $N_{t}$ for $U=18$, $\beta=1$, and $\kappa=1$.

### v2

@@ I Introduction
The Hybrid Monte Carlo (HMC) method is one of the most successful tools in the simulation of lattice field theories. However, despite its many advantages, its application often faces challenges due to manifolds of vanishing fermion determinant which result in infinite potential barriers that lead to diverging force terms in the molecular dynamics evolution. This separates regions in configuration space and results in ergodicity violations of the algorithm. One important example is the freezing of global topological fluctuations in lattice quantum chromodynamics (QCD) discretizations with chiral fermions [1] or near the continuum limit, where the integer-valued topological charge cannot be smoothly changed by molecular dynamics. Another prime example that exhibits configuration space partitioning is the fermionic Hubbard model in the particle/hole basis using the exponential discretization and formulated in terms of a non-compact auxiliary field by means of a continuous Hubbard-Stratonovich transformation. In this scenario, the action diverges due to vanishing fermion determinants, giving rise to insurmountable potential barriers that trigger an ergodicity problem in HMC simulations [2]. Reliable application of HMC necessitates the development of strategies to circumvent the potential barriers; Ref. [2] explores coarser molecular dynamics integration and discrete jumps while other approaches introduce a complexified auxiliary field over which to integrate [3, 4, 5, 6]. Normalizing flows represent a more radical departure from HMC for sampling these models [7] and do not use molecular dynamics evolution to generate the proposed updates. In this work, we propose another method, which interleaves the HMC simulation with so-called radial updates to overcome the ergodicity problem posed by the infinite potential barriers. Radial updates refer to multiplicative Metropolis-Hastings updates in a radial direction of non-compact fields that enable jumps over potential barriers and thus ensure ergodicity at comparably low computational cost.
@@ II Hybrid Monte Carlo
Critical to the correctness of HMC is that the combination of refreshing the momentum and MD integration must be able to reach every configuration of $\phi$ of nonzero weight. If infinite potential barriers partition the configuration space, the MD trajectory will never reach some valid configurations, leading to an unwanted bias in observable estimates. Notice that when $S$ diverges so too does the force $F[\phi]=-\frac{\partial S}{\partial\phi}$ (6) and the MD trajectory is rightly repelled. But if $S$ diverges everywhere on a (d1)(d-1)-dimensional manifold in the $d$-dimensional configuration space, i.e. on a codimension-1 manifold, the MD evolution becomes trapped and cannot reach every configuration. This bias is called an ergodicity problem.
In a practical implementation of HMC, the integration of Hamilton’s equations is inevitably done numerically. To this end, a trajectory of length $T$ is discretized into $N_{\mathrm{MD}}$ equal steps of size $\epsilon=T/N_{\mathrm{MD}}$. An initial sample $(\phi,\pi)\equiv(\phi(0),\pi(0))$ is then evolved to a new sample $(\tilde{\phi},\tilde{\pi})\equiv(\phi(T),\pi(T))$ using a numerical integration scheme. While an area-preserving and reversible integrator ensures the phase space is conserved through the integration, discretization errors can introduce violations of the energy conservation. Therefore, to retain detailed balance, the new configuration $\tilde{\phi}$ is proposed to the Markov chain and accepted according to the Metropolis test with probability,
@@ IV.1 Formulation
Here, $h=\Delta_{t}\delta_{\langle z,z^{\prime}\rangle}$ is the hopping matrix and $\mathcal{B}_{t}$ encodes the anti-periodic boundary conditions with $\mathcal{B}_{t}=+1$ for $0<t^{\prime}<N_{t}$ and $\mathcal{B}_{0}=-1$. This exponential discretization of the Hubbard model will serve as the basis for the analysis carried out in the remainder of this work. Other discretizations exist, but the exponential discretization is attractive because it preserves an exact chiral symmetry of the Hubbard Hamiltonian (12). However, that same symmetry implies that the fermion determinant $\det\left(M[\phi|\kappa]M[-\phi|-\kappa]\right)$ has manifolds of zero weight that divide the configuration space into distinct sectors [2]. As discussed in Section II, this makes HMC susceptible to ergodicity problems caused by infinite potential barriers that need to be addressed to ensure correct sampling.
@@ IV.2 Two-site model: restoring ergodicity
which eliminates autocorrelations arising from the harmonic part of the action [21]. The number of MD steps is tuned to obtain a fine integrator with acceptance rate $99\%$, leading to $N_{\mathrm{MD}}=60$ for $N_{x}=2$, $N_{t}=1$. We begin by generating $10^{4}$ configurations, saving the configuration after each HMC step. In the simulation with radial updates, we additionally employ one radial update per HMC step with standard deviation $\sigma_{R}=1.8$.
@@
Figure 2: HMC configurations (red dots) for the two-site model on a single time slice ($N_{t}=1$) for $U=18$, $\beta=1$ and $\kappa=1$. The configurations are compared to the exact distribution, shown as a contour plot. The HMC simulations were conducted using (17) and $N_{\mathrm{MD}}=60$, achieving an acceptance rate $99\%$. The left panel shows the simulation without radial updates, while the right panel includes radial updates. In the latter, a single radial update was performed before each HMC step with a proposal standard deviation of $\sigma_{R}=1.8$. Each panel shows $10^{4}$ recorded configurations, with measurements taken after each HMC step. The plots at the margins compare the exact marginal distribution to the histograms obtained from the visualized trajectories.
@@
Next, we increase the dimensionality $d$ of the system by increasing the number of time slices $N_{t}$, while keeping the parameters $U$, $\beta$, and $\kappa$ unchanged. The trajectory length $T$ is adjusted to reduce autocorrelations (17) and $N_{\mathrm{MD}}$ is tuned to maintain a fine integrator with an acceptance rate $99\%$. The specific parameter choices and further details on the simulations are given in Appendix B.2. In this case, the potential barriers can still be visualized by considering $\Phi_{x}=\sum_{t}\phi_{tx}$, because in the strong-coupling limit $U/\kappa\gg 1$, the probability weights of a configuration $\Phi=(\Phi_{1},\Phi_{2})$ are well-approximated by the exact one-site distribution derived in Ref. [2]. We begin by examining the example case of $N_{t}=8$, visualizing $10^{4}$ recorded configurations, both with and without radial updates. In both simulations, we recorded a configuration only after every 10th step, such that the shown trajectory corresponds to a total of $10^{5}$ iterations of the respective algorithm. In the simulation with radial updates we employ one radial update with standard deviation $\sigma_{R}=0.6$ per HMC step. The results are shown in Fig. 3. Similarly to the $N_{t}=1$ case, we observe that without radial updates, the trajectory remains confined to the middle diagonal band in the two-dimensional $\Phi$-plane. However, with radial updates, the algorithm efficiently explores the entire configuration space, effectively restoring ergodicity.
@@
Figure 3: HMC configurations (red dots) for the two-site model with $N_{t}=8$ for $U=18$, $\beta=1$ and $\kappa=1$. The configurations $\Phi=(\Phi_{1},\Phi_{2})$ with $\Phi_{x}=\sum_{t}\phi_{tx}$ are compared to the exact one-site distribution, recovered in the strong-coupling limit $U/\kappa\gg 1$ and shown as a contour plot. The HMC simulations were conducted using an optimal trajectory length (17) and $N_{\mathrm{MD}}=60$, achieving an acceptance rate $99\%$. The left panel shows the simulation without radial updates, while the right panel includes radial updates. In the latter, a single radial update was performed before each HMC step with a proposal standard deviation of $\sigma_{R}=0.6$. Each panel shows $10^{4}$ recorded configurations, with measurements taken after every 10th HMC step. The plots on the margins compare the marginal one-site distribution to the histograms obtained from the visualized trajectories.
@@ IV.3 Two-site model: reducing autocorrelation times
Figure 5: Integrated autocorrelation time $\tau_{\mathrm{int}}$ as a function of the proposal standard deviation $\sigma_{R}$ for the (left) observable $\mathcal{O}_{\Phi}$ (23) and the (right) single particle correlator $\mathcal{O}_{C}$ (19). The fits (green line) are obtained using the ansatz (25), with fit results shown in the corresponding legends. These results are used to estimate the optimal proposal standard deviation $\sigma_{R}^{\mathrm{(min)}}$ (grey dashed line), whose value is also provided in the legends. The underlying model is the two-site model with $N_{t}=8$ for $U=18$, $\beta=1$, and $\kappa=1$. The HMC simulations were conducted using an optimal trajectory length (17) and $N_{\mathrm{MD}}=50$, achieving an acceptance rate $99\%$. Each HMC step was preceded by a single radial update and a total of $N_{\mathrm{conf}}=3\times 10^{5}$ configurations were recorded, with measurements taken after each HMC step.
@@
Figure 6: The left panel shows a double logarithmic plot of the position of the minimum $\sigma_{R}^{(\mathrm{min})}$ as a function of dimensionality $d=N_{x}N_{t}$. The right panel illustrates a double logarithmic plot of the integrated autocorrelation time at the respective minimum, denoted by $\tau_{\mathrm{int}}^{(\mathrm{min})}$, also as a function of the dimensionality $d$. The data in both panels is fitted to the respective ansatz, (26) and (27), to determine the leading-order scaling. The results were obtained using $\mathcal{O}_{\Phi}$ (23) shown as blue circles, and the single particle correlator $\mathcal{O}_{C}$ (19) projected to the negative-parity irrep (20) shown as red squares. Data points shown in low opacity have been excluded from the fitting procedure and fit results are provided in the legends, where $\alpha^{\prime}=\log_{\mathrm{10}}\alpha$. The underlying model is the two-site Hubbard model with varying $N_{t}$ for $U=18$, $\beta=1$, and $\kappa=1$.
@@ IV.4 Perylene
As a final step in our analysis we involve radial updates in a realistic simulation of the molecule $C_{20}H_{12}$ perylene at half-filling. The molecule has raised interest due to its applications in organic electronics, such as solar cells [25, 26, 27] and semiconductors [28, 29], and was recently subject to a Hubbard-model study using HMC in Ref. [30]. In close similarity to this study, we set $\kappa=1$, $\beta=4,8$, $U=2$, and $N_{t}=96$. [^2] For reliable physical insight it is essential to perform the continuum extrapolation $N_{t}\rightarrow\infty$. For our purposes, it suffices to examine the algorithm’s performance at sufficiently large $N_{t}$. We employ an HMC with trajectory length (17) and $N_{\mathrm{MD}}=6,10$, resulting in a coarse MD integration with acceptance rates of about $70\%$. Before every HMC step, we employ a single radial update and we record a total of $N_{\mathrm{conf}}=2\times 10^{6}$ configurations, saving after each HMC step. As in Sec. IV.3, we perform multiple simulations across a range of $\sigma_{R}$ values to examine the tuning of the radial updates and to assess their impact on autocorrelation times. Specifically, we compute the integrated autocorrelation times for two representative observables. The observables considered are the $L_{0}$-norm of the field, defined as
@@ V Summary and Outlook
Using the Hubbard model as a primary example, we demonstrated that radial updates successfully restore ergodicity by enabling transitions between regions separated by infinitely high potential barriers caused by vanishing fermion determinants. Furthermore, we examined the scaling behavior of integrated autocorrelation times with the inclusion of radial updates, consistently observing polynomial scaling, which also indicates the absence of ergodicity violations in practice.
@@ B.2 Two-site model
For the two-site model, we performed a series of simulations for an increasing number of time slices ($N_{t}$), utilizing radial updates with varying proposal standard deviations ($\sigma_{R}$). For each choice of $N_{t}$, we employed an optimal HMC trajectory length (17) and tuned the number of MD steps ($N_{\mathrm{MD}}$) to obtain an acceptance rate of $99\%$. One radial update was applied prior to each HMC step, and measurements were taken after every HMC step. To obtain reliable estimates of the integrated autocorrelation times, the number of recorded configurations ($N_{\mathrm{conf}}$) was increased incrementally for higher dimensionalities. The specific choices of $N_{t}$, the corresponding $N_{\mathrm{MD}}$, and $N_{\mathrm{conf}}$ used in these simulations are summarized in Tab. 2.
@@ Appendix C Four-site model
Figure 12: Integrated autocorrelation time $\tau_{\mathrm{int}}$ as a function of the proposal standard deviation $\sigma_{R}$ for the observables $\mathcal{O}_{\Phi}$ (left) and $\mathcal{O}_{f}$ (right). The observables are defined in (23) and (30), respectively. The fits (green line) are obtained using the ansatz (25), with fit results shown in the corresponding legends. These results are used to estimate the optimal proposal standard deviation $\sigma_{R}^{\mathrm{(min)}}$ (grey dashed line), whose value is also provided in the legends. The underlying model is the four-site Hubbard model on $2\times 2$ square lattice with $N_{t}=8$ for $U=18$, $\beta=1$, and $\kappa=1$. The HMC simulations were conducted using (17) and $N_{\mathrm{MD}}=50$, achieving an acceptance rate $99\%$. Each HMC step was preceded by a single radial update and a total of $N_{\mathrm{conf}}=3\times 10^{5}$ configurations were recorded, with measurements taken after each HMC step.
Figure 13: The left panel shows a double logarithmic plot of the position of the minimum $\sigma_{R}^{(\mathrm{min})}$ as a function of dimensionality $d=N_{x}N_{t}$. The right panel illustrates a double logarithmic plot of the integrated autocorrelation time at the respective minimum, denoted by $\tau_{\mathrm{int}}^{(\mathrm{min})}$, also as a function of the dimensionality $d$. The data in both panels is fitted to the respective ansatz, (26) and (27), to determine the leading-order scaling. Fit results are provided in the legends, where $\alpha^{\prime}=\log_{\mathrm{10}}\alpha$. The underlying model is the four-site Hubbard model on a $2\times 2$ square lattice with varying $N_{t}$ for $U=18$, $\beta=1$, and $\kappa=1$.